A government has recently stressed its commitment to the future of youth and accordingly will subsidize you $2 for every $1 you spend on textbooks up to a max of $2100. Your income is $1500 and the price of books is $30 and the price of all other goods is $10. Whats the intercept of the original budget line? What’s the intercept for the budget line if there’s a grant with no maximum? At what point would you obtain max grant? What is the amount of conditional grant?
1
Expert's answer
2020-04-27T07:37:35-0400

1.X=50, Y=150

"\frac{1500}{30}=50" "\frac{1500}{10}=150"

2.X=150, Y=150

"\frac{4500}{30}=150" 1500+3000=4500

3000 - subsidy

3.X=105, Y=45

3600=30x+10y, 3600=2100+1500

4.2100